# Small Language Models are the Future of Agentic AI

### v1

@@ Abstract
Our position, formulated as a value statement, highlights the significance of the operational and economic impact even a partial shift from LLMs to SLMs is to have on the AI agent industry. We aim to stimulate the discussion on the effective use of AI resources and hope to advance the efforts to lower the costs of AI of the present day. Calling for both contributions to and critique of our position, we commit to publishing all such correspondence at research.nvidia.com/labs/lpr/slm-agents.
@@ 1 Introduction
LLM API endpoints are specifically designed to serve a large volume of diverse requests using one generalist LLM. This operational model is deeply ingrained in the industry — so deeply ingrained, in fact, that it forms the foundation of substantial capital bets: While the market for the LLM API serving that underlies agentic applications was estimated at USD 5.6bn in 2024 grandview2025llm , the investment into the hosting cloud infrastructure surged to USD 57bn in the same year yadav2025ai . The 10-fold discrepancy between investment and market size has been accepted, because it is assumed that this operational model will remain the cornerstone of the industry without any substantial alterations, and that the large initial investment will deliver returns comparable to traditional software and internet solutions within 3-4 years morganstanley2025genai .
In this work, we recognize the dominance of the standard operational model but verbally challenge one of its aspects, namely the custom that the agents’ requests to access language intelligence are – in spite of their comparative simplicity – handled by singleton choices of generalist LLMs. We state (Section 2), argue (Section 3), and defend (Section 4) the position that the small, rather than large, language models are the future of agentic AI. We, however, recognize the business commitment and the now-legacy praxis that is the cause for the contrary state of the present (Section 5). In remedy, we provide an outline of a conversion algorithm for the migration of agentic applications from LLMs to SLMs (Section 6), and call for a wider discussion (Section 7). If needed to concretize our stance, we attach a set of short case studies estimating the potential extent of LLM-to-SLM replacement in selected popular open-source agents (Appendix B).
@@ 2. WD2
We justify the wording of these definitions in Appendix A, but note that their choice has little bearing on the essence of our position. We note that as of 2025, we would be comfortable with considering most models below 10bn parameters in size to be SLMs.
@@ 2.2 Statement
and that on the basis of items V1, V2 and V3 SLMs are the future of agentic AI.
@@ 3 Position Arguments
We support items V1, V2 and V3 by the following non-exclusive arguments.
@@ 3.1 SLMs are already sufficiently powerful for use in agents
    SLMs are sufficiently powerful to take the place of LLMs in agentic systems. This argument stands in support of item V1.
@@
Extensive comparisons with large models have been conducted in the individual works cited below, but not all capabilities assessed by benchmarks are essential to their deployment in the agentic context. Here we highlight their aptitude for commonsense reasoning (an indicator of basic understanding), tool calling and code generation (both indicators of the ability to correctly communicate across the model→tool/code interface; see Figure 1; berkeley-function-calling-leaderboard ; yao2024tau ), and instruction following (ability to correctly respond back across the code←model interface; zhou2023instruction ). In each case, we also quote the efficiency increase if stated by the authors.
  • Microsoft Phi series. Phi-2 (2.7bn) achieves commonsense reasoning scores and code generation scores on par with 30bn models while running $\sim$15$\times$ faster phi22023 . Phi-3 small (7bn) abdin2024phi achieves language understanding and commonsense reasoning on par with and code generation scores running up to 70bn models of the same generation.
@@
  • NVIDIA Hymba-1.5B. This Mamba-attention hybrid-head SLM demonstrates best instruction accuracy and 3.5$\times$ greater token throughput than comparably-sized transformer models dong2024hymba . On instruction following, it outperforms larger 13bn models.
@@
Note that on top of competitive off-the-shelf performance, the reasoning capabilities of SLMs can be enhanced at inference time with self-consistency, verifier feedback, or tool augmentation — e.g., Toolformer (6.7bn) outperforms GPT-3 (175bn) via API use toolformer2023 , and 1-3bn models have rivaled 30bn+ LLMs on math problems via structured reasoning zhou2022leasttomost .
@@ 3.2 SLMs are more economical in agentic systems
    SLMs are more economical than LLMs in agentic systems. This argument supports item V3.
Small models provide significant benefits in cost-efficiency, adaptability, and deployment flexibility. These advantages are specifically valuable in agentic workflows where specialization and iterative refinement are critical. Section 3.1 detailed a number of efficiency comparisons of the listed SLMs to relevant LLMs. Here we draw a more encompassing picture to support item A2.
  • Inference efficiency. Serving a 7bn SLM is 10–30$\times$ cheaper (in latency, energy consumption, and FLOPs) than a 70–175bn LLM, enabling real-time agentic responses at scale subramanian2025small ; shone2024explore ; invisible2025slm ; mehta2024energy . Recent advances in inference operating systems such as NVIDIA Dynamo dynamo2025 explicitly provide support for high-throughput, low-latency SLM inference in both cloud and edge deployments. In addition, since SLMs require less or no parallelization across GPUs and nodes, the maintenance and operation of the serving infrastructure comes at a lower expense as well (see item CA4 and item A13).
  • Fine-tuning agility. Parameter-efficient (e.g., LoRA hu2021lora and DoRA liu2024dora ) and full-parameter finetuning for SLMs require only a few GPU-hours, allowing behaviors to be added, fixed, or specialized overnight rather than over weeks subramanian2025small .
@@
  • Parameter utilization. At the outset, LLMs appear to operate as monoliths involving a large amount of parameters representing swathes of compressed information in the production of their outputs. On a closer look, however, much of the signals passing through these systems is sparse, engaging only a fraction of their parameters for any single input song2024powerinfer ; liu2023deja . That this behavior appears to be more subdued in SLMs song2024powerinfer ; xue2024powerinfer suggests that SLMs may be fundamentally more efficient by the virtue of having a smaller proportion of their parameters contribute to the inference cost without a tangible effect on the output.
@@ 3.3 SLMs are more flexible
    SLMs possess greater operational flexibility in comparison to LLMs. This argument stands in support of items V2 and V3.
Due to their small size and the associated reduction in pre-training and fine-tuning costs (Section 3.2), SLMs are inherently more flexible than their large counterparts when appearing in agentic systems. As such, it becomes much more affordable and practical to train, adapt, and deploy multiple specialized expert models for different agentic routines. This efficiency enables rapid iteration and adaptation, making it feasible to address evolving user needs, including supporting new behaviors, meeting new output formatting requirements, and complying with changing local regulation in selected markets wang2024comprehensive ; kumar2025large ; thamm2024trustworthy .
@@ 3.4 Agents expose only very narrow LM functionality
    Agentic applications are interfaces to a limited subset of LM capabilities. This supports items V1 and V2.
@@
It could be argued back that the careful interfacing with a generalist LLM is necessary for strong performance on the narrow task because of the LLM’s better understanding of the broader language and the world (item AV1). This is addressed in Section 4.1.
@@ 3.5 Agentic interactions necessitate close behavioral alignment
    Agentic interactions necessitate close behavioral alignment. This aligns with item V2.
@@ 3.6 Agentic systems are naturally heterogeneous
    Agentic systems naturally allow for heterogeneity in the selection of models that they use. This aligns with item V2.
A language model can itself be a tool called by another language model. Likewise, every time the agent’s code invokes a language model, it can, in principle, choose any language model. This is illustrated in Figure 1. We argue that incorporating multiple language models of different sizes and capabilities for queries or operations of different levels of complexity offers a natural way for the introduction of SLMs. In the context of Figure 1-Left, an LLM can be used for the model with the root agency, while a SLM could be used for the subordinate LM. In Figure 1-Right, all LMs could in principle be specialized SLMs: one for conversationality, another one for carrying out controller-defined language modeling tasks.
@@ 3.7 Agentic interactions are natural pathways for gathering data for future improvement
    Agentic interactions are a good source for data for future model improvement. This is fundamentally supportive of item V2.
As noted in Section 3.4, invocations of tools and language models during an agentic process are often accompanied by careful prompting that focuses the language model on delivering the narrow functionality that is required at the time. Each one of these invocations is itself a natural source of data for future improvement (under the necessary assumption that no non-retainable confidential data is being processed). A listener decorating the tool/model call interface can gather specialized instruction data that can later be used to produce a fine-tune an expert SLM and lower the cost of that call in the future (see logger in Figure 1). We argue that this avenue is enabled by the architecture of agents masterman2024landscape and produces high-quality organic data (that can be further post-filtered by considering the overall success of the workflow), thus making the production of expert SLMs to stand in place of LLMs a natural step in agent deployment — not just an auxiliary effort.
@@ 1. AV1
This alternative view disputes item V2 and rests on the following counter-arguments:
@@ 2. CA2
We believe that item CA1 is too limited to attack item V2, namely because
@@
    The flexibility of small language models (Section 3.3) comes to the rescue. A small language model can be easily fine-tuned for the task $\mathcal{T}$ of item AV1 to perform to the desired level of reliability. This is unaccounted for in scaling law studies.
@@ 3. A10
    Reasoning (or, more generally, test-time compute scaling; see Section 3.2) is significantly more affordable. A small language model, still retaining its benefits of greater cross-device agility can be reasonable expected to be scalable at inference time to the desired level of reliability.
We also believe that item CA2 is too arcane to attack item V2 because
@@ 1. AV2
It could be argued that the analysis in item A2 that put forth in favor of item V3 was ignorant of the wider business of AI model deployment:
@@ 2. CA4
Acknowledgment. We acknowledge that item AV2 is a valid view, with the exact economical considerations being highly case-specific. We believe that the jury is still out on item AV2, but that several factors hint that item V3 might prevail:
@@ 1. A12
    Recent improvements in inference scheduling and large inference system modularization offer unprecedented levels of inference system flexibility in monolithic computing clusters zier2025dynamo ; ai-dynamo-dynamo ; mann2025dynamo , countering the traditional stance expressed in item CA3.
@@ 1. AV3
Acknowledgment. We acknowledge item AV3 as a distinct possibility, but maintain the position that the weight of advantages described across items A1, A2, A3, A4, A5, A6 and A7 can plausibily overturn the present state of affairs.
@@ 5 Barriers to Adoption
It would be prudent to ask oneself: If the items A1, A2, A3, A4, A5, A6 and A7 are truly compelling, why do the ever newer generations of agents seemingly just perpetuate the status quo of using generalist LLMs?
@@
    Large amounts of upfront investment into centralized LLM inference infrastructure. As detailed in Section 1, large capital bets have been made on the centralized LLM inference being the leading paradigm in providing AI services in the future. As such, the industry has been much quicker at building the tools and infrastructure to that end, omitting any considerations for the possibility that more decentralized SLM or on-device inference might be equally feasible in the near future.
@@
We note that items B1, B2 and B3 are practical hurdles and far from being fundamental flaws of the SLM technology in the context of agentic AI. With advanced inference scheduling systems such as Dynamo dynamo2025 , item B1 is being reduced to a mere effect of inertia. item B2 is becoming increasingly recognized in the field dong2024hymba ; phi22023 , and it would be natural for item B3 to fall once the economic benefits of SLM deployment in agentic applications (item A2) are better known. With the inertia of item B1 in particular, we do not endeavor to give a timeline for the retreat of these barriers or the popular adoption of SLMs.
@@ 6 LLM-to-SLM Agent Conversion Algorithm
    Secure usage data collection. The initial step involves deploying instrumentation to log all non-HCI agent calls, capturing input prompts, output responses, contents of individual tool calls, and optionally latency metrics for a later targeted optimization. In terms of implementation, it is the recommended practice to set up encrypted logging pipelines with role-based access controls metomic2025aiagents and anonymize all data with respect to its origins before storage workos2025secureagents . See logger in Figure 1 for an illustration.
@@
    Data curation and filtering. Begin collecting data through the pipelines of item S1. Once a satisfactory amount of data has been collected (10k-100k examples being sufficient for fine-tuning of small models as a rule of thumb agarwal2024delift ; ding2023parameter ), it is necessary to remove any PII, PHI, or any other application-specific sensitive data that could cause a data leak across user accounts once used to produce a SLM specialist. Many typical varieties of sensitive data can be detected and masked or removed using popular automated tools for dataset preparation sainz2024open ; radhakrishnan2023certified . Application specific inputs (e.g. legal or internal documents) can be often be automatically paraphrased to obfuscate named entities and numerical details without compromising the general information content brennan2012adversarial ; yu2024privacy ; yan2024protecting .
@@
    SLM selection. For each identified task, select one or more candidate SLMs. Criteria for selection include the SLM’s inherent capabilities (e.g., instruction following, reasoning, context window size), its performance on relevant benchmarks for the task type, its licensing, and its deployment footprint (memory, computational requirements). Models of Section 3.2 serve as good starting candidates.
@@
    Specialized SLM fine-tuning. For each selected task and corresponding SLM candidate, prepare a task-specific dataset from the curated data collected in items S2 and S3. Then, fine-tune the chosen SLMs on these specialized datasets. PEFT techniques such as LoRA hu2022lora or QLoRA dettmers2023qlora can be leveraged to reduce computational costs and memory requirements associated with fine-tuning, making the process more accessible. Full fine-tuning can also be considered if resources permit and maximal adaptation is required. In some cases, it may be beneficial to use knowledge distillation, where the specialist SLM is trained to mimic the outputs of the more powerful generalist LLM on the task-specific dataset. This can help transfer some of the more nuanced capabilities of the LLM to the SLM.
@@
    Iteration and refinement. One may retrain the SLMs and the router model periodically with new data to maintain performance and adapt to evolving usage patterns. This forms a continuous improvement loop, returning to item S2 or item S4 as appropriate.

### v2

@@ Abstract
Our position [^1] The views and positions expressed in this paper are those of the authors and do not necessarily reflect the views or positions of any entities they represent., formulated as a value statement, highlights the significance of the operational and economic impact even a partial shift from LLMs to SLMs is to have on the AI agent industry. We aim to stimulate the discussion on the effective use of AI resources and hope to advance the efforts to lower the costs of AI of the present day. Calling for both contributions to and critique of our position, we commit to publishing all such correspondence at research.nvidia.com/labs/lpr/slm-agents.
@@ 1 Introduction
LLM API endpoints are specifically designed to serve a large volume of diverse requests using one generalist LLM. This operational model is deeply ingrained in the industry. In fact, it forms the foundation of substantial capital investment in the hosting cloud infrastructure – estimated at USD 57bn colliers2025datacenter . It is anticipated that this operational model will remain the cornerstone of the industry and that the large initial investment will deliver returns comparable to traditional software and internet solutions within 3-4 years morganstanley2025genai .
In this work, we recognize the dominance of the standard operational model but verbally challenge one of its aspects, namely the custom that the agents’ requests to access language intelligence are – in spite of their comparative simplicity – handled by singleton choices of generalist LLMs. We state (Section˜2), argue (Section˜3), and defend (Section˜4) the position that the small, rather than large, language models are the future of agentic AI. We, however, recognize the business commitment and the now-legacy praxis that is the cause for the contrary state of the present (Section˜5). In remedy, we provide an outline of a conversion algorithm for the migration of agentic applications from LLMs to SLMs (Section˜6), and call for a wider discussion (Section˜7). If needed to concretize our stance, we attach a set of short case studies estimating the potential extent of LLM-to-SLM replacement in selected popular open-source agents (Appendix˜B).
@@ 2. WD2
We justify the wording of these definitions in Appendix˜A, but note that their choice has little bearing on the essence of our position. We note that as of 2025, we would be comfortable with considering most models below 10bn parameters in size to be SLMs.
@@ 2.2 Statement
and that on the basis of items˜V1, V2 and V3 SLMs are the future of agentic AI.
@@ 3 Position Arguments
We support items˜V1, V2 and V3 by the following non-exclusive arguments.
@@ 3.1 SLMs are already sufficiently powerful for use in agents
    SLMs are sufficiently powerful to take the place of LLMs in agentic systems. This argument stands in support of item˜V1.
@@
Extensive comparisons with large models have been conducted in the individual works cited below, but not all capabilities assessed by benchmarks are essential to their deployment in the agentic context. Here we highlight their aptitude for commonsense reasoning (an indicator of basic understanding), tool calling and code generation (both indicators of the ability to correctly communicate across the model→tool/code interface; see Figure˜1; berkeley-function-calling-leaderboard ; yao2024tau ), and instruction following (ability to correctly respond back across the code←model interface; zhou2023instruction ). In each case, we also quote the efficiency increase if stated by the authors.
[…]
@@
[…]
@@
Note that on top of competitive off-the-shelf performance, the reasoning capabilities of SLMs can be enhanced by light-weight selective finetuning lore2024large or at inference time with self-consistency, verifier feedback, or tool augmentation — e.g., Toolformer (6.7bn) outperforms GPT-3 (175bn) via API use toolformer2023 , and 1-3bn models have rivaled 30bn+ LLMs on math problems via structured reasoning zhou2022leasttomost .
@@ 3.2 SLMs are more economical in agentic systems
    SLMs are more economical than LLMs in agentic systems. This argument supports item˜V3.
Small models provide significant benefits in cost-efficiency, adaptability, and deployment flexibility. These advantages are specifically valuable in agentic workflows where specialization and iterative refinement are critical. Section˜3.1 detailed a number of efficiency comparisons of the listed SLMs to relevant LLMs. Here we draw a more encompassing picture to support item˜A2.
[…]
  • Fine-tuning agility. Parameter-efficient (e.g., LoRA hu2021lora and DoRA liu2024dora ), low-resource belcak2025minifinetuning , and full-parameter finetuning for SLMs require only a few GPU-hours, allowing behaviors to be added, fixed, or specialized overnight rather than over weeks subramanian2025small .
@@
  • Parameter and embedding space utilization. At the outset, LLMs appear to operate as monoliths involving a large amount of parameters representing swathes of compressed information in the production of their outputs. On a closer look, however, many of the embeddings passing through these systems are very sparse, engaging only a fraction of their parameters for any single input song2024powerinfer ; liu2023deja and being effectively compressible belcak2024tiny ; gu2025text . That this behavior appears to be more subdued in SLMs song2024powerinfer ; xue2024powerinfer suggests that SLMs may be fundamentally more efficient by the virtue of having a smaller proportion of their parameters contribute to the inference cost without a tangible effect on the output.
@@ 3.3 SLMs are more flexible
    SLMs possess greater operational flexibility in comparison to LLMs. This argument stands in support of items˜V2 and V3.
Due to their small size and the associated reduction in pre-training and fine-tuning costs (Section˜3.2), SLMs are inherently more flexible than their large counterparts when appearing in agentic systems. As such, it becomes much more affordable and practical to train, adapt, and deploy multiple specialized expert models for different agentic routines. This efficiency enables rapid iteration and adaptation, making it feasible to address evolving user needs, including supporting new behaviors, meeting new output formatting requirements, and complying with changing local regulation in selected markets wang2024comprehensive ; kumar2025large ; thamm2024trustworthy .
@@ 3.4 Agents expose only very narrow LM functionality
    Agentic applications are interfaces to a limited subset of LM capabilities. This supports items˜V1 and V2.
@@
It could be argued back that the careful interfacing with a generalist LLM is necessary for strong performance on the narrow task because of the LLM’s better understanding of the broader language and the world (item˜AV1). This is addressed in Section˜4.1.
@@ 3.5 Agentic interactions necessitate close behavioral alignment
    Agentic interactions necessitate close behavioral alignment. This aligns with item˜V2.
@@ 3.6 Agentic systems are naturally heterogeneous
    Agentic systems naturally allow for heterogeneity in the selection of models that they use. This aligns with item˜V2.
A language model can itself be a tool called by another language model. Likewise, every time the agent’s code invokes a language model, it can, in principle, choose any language model. This is illustrated in Figure˜1. We argue that incorporating multiple language models of different sizes and capabilities for queries or operations of different levels of complexity offers a natural way for the introduction of SLMs. In the context of Figure˜1-Left, an LLM can be used for the model with the root agency, while a SLM could be used for the subordinate LM. In Figure˜1-Right, all LMs could in principle be specialized SLMs: one for conversationality, another one for carrying out controller-defined language modeling tasks.
@@ 3.7 Agentic interactions are natural pathways for gathering data for future improvement
    Agentic interactions are a good source for data for future model improvement. This is fundamentally supportive of item˜V2.
As noted in Section˜3.4, invocations of tools and language models during an agentic process are often accompanied by careful prompting that focuses the language model on delivering the narrow functionality that is required at the time. Each one of these invocations is itself a natural source of data for future improvement (under the necessary assumption that no non-retainable confidential data is being processed). A listener decorating the tool/model call interface can gather specialized instruction data that can later be used to produce a fine-tune an expert SLM and lower the cost of that call in the future (see logger in Figure˜1). We argue that this avenue is enabled by the architecture of agents masterman2024landscape and produces high-quality organic data (that can be further post-filtered by considering the overall success of the workflow), thus making the production of expert SLMs to stand in place of LLMs a natural step in agent deployment — not just an auxiliary effort.
@@ 1. AV1
This alternative view disputes item˜V2 and rests on the following counter-arguments:
@@ 2. CA2
We believe that item˜CA1 is too limited to attack item˜V2, namely because
@@
    The flexibility of small language models (Section˜3.3) comes to the rescue. A small language model can be easily fine-tuned for the task $\mathcal{T}$ of item˜AV1 to perform to the desired level of reliability. This is unaccounted for in scaling law studies.
@@ 3. A10
    Reasoning (or, more generally, test-time compute scaling; see Section˜3.2) is significantly more affordable. A small language model, still retaining its benefits of greater cross-device agility can be reasonable expected to be scalable at inference time to the desired level of reliability.
We also believe that item˜CA2 is too arcane to attack item˜V2 because
@@ 1. AV2
It could be argued that the analysis in item˜A2 that put forth in favor of item˜V3 was ignorant of the wider business of AI model deployment:
@@ 2. CA4
Acknowledgment. We acknowledge that item˜AV2 is a valid view, with the exact economical considerations being highly case-specific. We believe that the jury is still out on item˜AV2, but that several factors hint that item˜V3 might prevail:
@@ 1. A12
    Recent improvements in inference scheduling and large inference system modularization offer unprecedented levels of inference system flexibility in monolithic computing clusters zier2025dynamo ; ai-dynamo-dynamo ; mann2025dynamo , countering the traditional stance expressed in item˜CA3.
@@ 1. AV3
Acknowledgment. We acknowledge item˜AV3 as a distinct possibility, but maintain the position that the weight of advantages described across items˜A1, A2, A3, A4, A5, A6 and A7 can plausibily overturn the present state of affairs.
@@ 5 Barriers to Adoption
It would be prudent to ask oneself: If the items˜A1, A2, A3, A4, A5, A6 and A7 are truly compelling, why do the ever newer generations of agents seemingly just perpetuate the status quo of using generalist LLMs?
@@
    Large amounts of upfront investment in centralized LLM inference infrastructure. As detailed in Section˜1, large capital bets have been made on the centralized LLM inference being the leading paradigm in providing AI services in the future. As such, the industry has been much quicker at building the tools and infrastructure to that end, omitting any considerations for the possibility that more decentralized SLM or on-device inference might be equally feasible in the near future.
@@
We note that items˜B1, B2 and B3 are practical hurdles and far from being fundamental flaws of the SLM technology in the context of agentic AI. With advanced inference scheduling systems such as Dynamo dynamo2025 , item˜B1 is being reduced to a mere effect of inertia. item˜B2 is becoming increasingly recognized in the field dong2024hymba ; phi22023 , and it would be natural for item˜B3 to fall once the economic benefits of SLM deployment in agentic applications (item˜A2) are better known. With the inertia of item˜B1 in particular, we do not endeavor to give a timeline for the retreat of these barriers or the popular adoption of SLMs.
@@ 6 LLM-to-SLM Agent Conversion Algorithm
    Secure usage data collection. The initial step involves deploying instrumentation to log all non-HCI agent calls, capturing input prompts, output responses, contents of individual tool calls, and optionally latency metrics for a later targeted optimization. In terms of implementation, it is the recommended practice to set up encrypted logging pipelines with role-based access controls metomic2025aiagents and anonymize all data with respect to its origins before storage workos2025secureagents . See logger in Figure˜1 for an illustration.
@@
    Data curation and filtering. Begin collecting data through the pipelines of item˜S1. Once a satisfactory amount of data has been collected (10k-100k examples being sufficient for fine-tuning of small models as a rule of thumb agarwal2024delift ; ding2023parameter ), it is necessary to remove any PII, PHI, or any other application-specific sensitive data that could cause a data leak across user accounts once used to produce a SLM specialist. Many typical varieties of sensitive data can be detected and masked or removed using popular automated tools for dataset preparation sainz2024open ; radhakrishnan2023certified . Application specific inputs (e.g. legal or internal documents) can be often be automatically paraphrased to obfuscate named entities and numerical details without compromising the general information content brennan2012adversarial ; yu2024privacy ; yan2024protecting .
@@
    SLM selection. For each identified task, select one or more candidate SLMs. Criteria for selection include the SLM’s inherent capabilities (e.g., instruction following, reasoning, context window size), its performance on relevant benchmarks for the task type, its licensing, and its deployment footprint (memory, computational requirements). Models of Section˜3.2 serve as good starting candidates.
@@
    Specialized SLM fine-tuning. For each selected task and corresponding SLM candidate, prepare a task-specific dataset from the curated data collected in items˜S2 and S3. Then, fine-tune the chosen SLMs on these specialized datasets. PEFT techniques such as LoRA hu2022lora or QLoRA dettmers2023qlora can be leveraged to reduce computational costs and memory requirements associated with fine-tuning, making the process more accessible. Full fine-tuning can also be considered if resources permit and maximal adaptation is required. In some cases, it may be beneficial to use knowledge distillation, where the specialist SLM is trained to mimic the outputs of the more powerful generalist LLM on the task-specific dataset. This can help transfer some of the more nuanced capabilities of the LLM to the SLM.
@@
    Iteration and refinement. One may retrain the SLMs and the router model periodically with new data to maintain performance and adapt to evolving usage patterns. This forms a continuous improvement loop, returning to item˜S2 or item˜S4 as appropriate.